two cars start at the same point going opposite derections one travels 6mph faster than the other. in 4 houts they are 184 mile apart how fast were each going

Found 2 solutions by jorel1380, nerdybill:
Answer by jorel1380(3719)

4(p+p+6)=184
8p+48=184
8p=136
p=17
p+6=23
The first car was going 17 mph, the second car went 23 mph..

Answer by nerdybill(7384)

two cars start at the same point going opposite derections one travels 6mph faster than the other. in 4 houts they are 184 mile apart how fast were each going
.
Let x = speed of slower car
then
x+6 = speed of faster car
.
4x + 4(x+6) = 186
dividing both sides by 4:
x + x+6 = 46.5
2x+6 = 46.5
2x = 40.5
x = 20.25 mph (slower car)
.
faster car:
x+6 = 20.25+6 = 26.25 mph